OpenStudy (hockeychick23):

Suppose that 1% of the students in a school have head lice and the test for head lice is accurate 75% of the time. What is the probability that a student in the school has head lice, given that the test came back positive?

OpenStudy (perl):

P( lice ) = .01 P( test positive | lice ) = .75 Find probability P ( lice | test positive)

OpenStudy (perl):

Let's derive a rule P(A | B) = P(A & B) / P(B) now P(A&B) = P(A) * P(B | A ) so lets substitute that P(A|B) = P(A) P(B|A) / P(B) P ( lice | test pos. ) = P(lice) * P ( test pos. | lice ) / P( test pos.)

OpenStudy (perl):

does that make sense so far

OpenStudy (hockeychick23):

yea, i got to that point but after that i keep getting .75 as an answer

OpenStudy (perl):

it might be easier to use a tree diagram

OpenStudy (perl):

it says the test is accurate 75% of the time there are two types of 'accuracy. if someone has lice then it tests positive 75% of the time if someone does not have lice then it tests negative 75% of the time

OpenStudy (perl):



OpenStudy (perl):